The temperature of a sample of fluorine gas is changed without changing its volume, causing a change in pressure from 46.0 kPa to 129.0 kPa. If the starting temperature was 225 °C, what is the final temperature in degrees Celsius?

Answer value


1
Expert's answer
2021-03-17T06:39:42-0400

P1= 46.0kPa

T1 = 225+273 = 498 k



P2=129.0kPa


P/ T1 = P2 / T2


"\frac{46}{498}= \frac {129}{T2}"


"T2= \frac{498×129}{46}=1396.57"


Tin degrees celsius= (1396.57 - 273) = 1123.57 degrees Celsius